OpenStudy (anonymous):

Another question Sarah earned some money taking care of a dog. She charged $8.50 for washing the dog and $7 for walking the dog. She spent $1.15 on shampoo to wash the dog. What was Sarah’s profit from taking care of the dog?

OpenStudy (anonymous):

total money charged 8.50+7 =15.50

OpenStudy (anonymous):

total money earned =15.50-1.15 =14.35 $

OpenStudy (anonymous):

excuse but in which class are u

OpenStudy (anonymous):

grade 4

OpenStudy (anonymous):

okay heres a question for you

OpenStudy (anonymous):

john buys a computer for 100$ and he sells it for 90$ then again buy it buy for 80$ ?calculate the total loss or profit ?

OpenStudy (anonymous):

180$

OpenStudy (anonymous):

wait what is the answer?

OpenStudy (anonymous):

oh ok I know already thanks

OpenStudy (anonymous):

total loss is 20 $